# Large Language Models are In-Context Molecule Learners

### v2

@@ 3.2.2 Sequence Reversal
3.3 In-Context Molecule Tuning
As shown in Figure 1, similar molecules typically share similar structures and chemical properties. Building on this principle, MolReGPT  Li et al. (2023) has demonstrated the effectiveness of in-context learning, which aims to prompt Large Language Models (LLMs) to learn from similar examples without the need for domain-specific pre-training and fine-tuning. However, MolReGPT heavily relies on the reasoning and in-context learning capabilities of LLMs, resulting in poor performance with relatively small language models. To address the deficits , we propose In-Context Molecule Tuning to fine-tune the parameters of LLMs, enabling them to learn from context examples and reason on the current input. Notably, In-Context Molecule Tuning can be easily adapted to any LLM, allowing even smaller language models to unlock their in-context molecule learning capability by learning the differences and similarities between molecule-caption pairs in the context.
@@ 4 Experiments
In this section, we aim to evaluate the effectiveness of ICMA. Firstly, we introduce the experimental settings. Then, we compare ICMA with selected baseline models on the ChEBI-20 dataset and further test ICMA on a smaller dataset, PubChem324k. Furthermore, we also study the factors that will affect the performance of IMCA, including different foundation models, retrieval algorithms, context settings, and model scales. Lastly, an ablation study is also conducted to justify the design of Post-retrieval Re-ranking components.
4.1 Experimental Settings
@@
We will first detail our experiment settings. All the hyper-parameters are illustrated in Appendix A Table 10. If not specifically stated, the cutoff length is set to 1024 and n_shot is set to 2 to control variable. Notably, for LLMs with over 7 billion parameters, we apply LoRA  Hu et al. (2021) to save the GPU memory and accelerate computation. Otherwise, we fine-tune the full model of LLMs. For the dataset, we apply two different molecule-caption translation datasets, ChEBI-20  Edwards et al. (2021) and PubChem324k  Liu et al. (2023b). The details of the datasets are shown in Table 1.
Dataset Train Validation Test
ChEBI-20 26,407 3,001 3,000
PubChem324k 12,000 1,000 2,000
Table 1: Details of the datasets, ChEBI-20 and PubChem324k. For PubChem324k, we follow the split in  Liu et al. (2023b), while ignoring the Pretrain fold.
For comparison, we select two different foundation LLMs as the backbones of ICMA, namely Galactica-125M  Taylor et al. (2022) and Mistral-7B-instruct-v0.2  Jiang et al. (2023). The former is a representative and smaller LLM that has been pre-trained on unstructured scientific corpora, which have been aware of the molecular knowledge from the pre-training stage, while the latter is a general LLM with 7 billion parameters, whose capabilities are comparable to GPT-3.5-turbo. Notably, in the factor analysis and ablation study parts, we select Galactica-125M for experiments to reduce the computational costs. Additionally, considering baseline models, we mainly select MolT5-base, MolT5-large from  Edwards et al. (2022) and MolReGPT (GPT-3.5-turbo and GPT-4-0314) from  Li et al. (2023) for comparison on the ChEBI-20 dataset.
4.2 Performance Comparison
[…]
MolT5-base Edwards et al. (2022) 0.540 0.457 0.634 0.485 0.578 0.569
MolReGPT (GPT-3.5-turbo) 0.565 0.482 0.623 0.450 0.543 0.585
MolT5-large Edwards et al. (2022) 0.594 0.508 0.654 0.510 0.594 0.614
MolReGPT (GPT-4-0314) 0.607 0.525 0.634 0.476 0.562 0.610
Galactica-125M 0.585 0.501 0.630 0.474 0.568 0.591
ICMA(Galactica-125M)2,2048 0.636 0.565 0.674 0.536 0.615 0.648
Mistral-7B 0.566 0.478 0.614 0.449 0.547 0.572
ICMA(Mistral-7B)2,2048 0.651 0.581 0.686 0.550 0.625 0.661
Table 2: Mol2Cap results on ChEBI-20 dataset (Best, Second Best). Here, the results of MolT5-base and MolT5-large are domain-specific pre-training & fine-tuning results retrieved from Edwards et al. (2022), while MolReGPT(GPT-3.5-turbo) and MolReGPT(GPT-4-0314) are prompting & in-context learning results retrieved from Li et al. (2023). More importantly, Galactica-125M and Mistral-7B demonstrate vanilla supervised fine-tuned results, while ICMA(Galactica-125M) and ICMA(Mistral-7B) illustrate the ICMA results.
[…]
MolT5-base Edwards et al. (2022) 0.769 0.081 24.458 0.721 0.588 0.529 0.772
MolReGPT(GPT-3.5-turbo) 0.790 0.139 24.91 0.847 0.708 0.624 0.887
MolT5-large Edwards et al. (2022) 0.854 0.311 16.071 0.834 0.746 0.684 0.905
MolReGPT(GPT-4-0314) 0.857 0.280 17.14 0.903 0.805 0.739 0.899
Galactica-125M 0.781 0.173 26.34 0.836 0.708 0.631 0.916
ICMA(Galactica-125M)4,2048 0.836 0.342 21.48 0.893 0.809 0.743 0.934
Mistral-7B 0.767 0.234 27.39 0.852 0.718 0.649 0.918
ICMA(Mistral-7B)4,2048 0.855 0.460 18.73 0.916 0.837 0.789 0.958
Table 3: Cap2Mol results on ChEBI-20 dataset (Best, Second Best). Here, the results of MolT5-base and MolT5-large are domain-specific pre-training & fine-tuning results retrieved from Edwards et al. (2022), while MolReGPT(GPT-3.5-turbo) and MolReGPT(GPT-4-0314) are prompting & in-context learning results retrieved from Li et al. (2023). More importantly, Galactica-125M and Mistral-7B demonstrate vanilla supervised fine-tuned results, while ICMA(Galactica-125M) and ICMA(Mistral-7B) illustrate the ICMA results.
Mol2Cap. As illustrated in Table 2, Galactica-125M with SFT has already shown competitive performance to previous baselines due to its pre-training on scientific corpora. However, ICMA could still improve the performance of Galactica-125M by 12.8% and 8.3% considering the BLEU-4 and ROUGE-L scores on ChEBI-20 dataset. With only 125 million parameters, ICMA(Galactica-125M) can beat MolT5-large, which owns 780 million parameters. Meanwhile, the general LLM, Mistral-7B with naive SFT, only achieves a performance that is slightly better than MolT5-base, despite the fact that Mistral-7B is 70 times larger than MolT5-base. This outcome is not surprising because Mistral-7B is not specifically designed or pre-trained for biomolecular purposes. It also reveals that although general LLMs have illustrated powerful capabilities with billions of parameters, few works have adapted them to the biomolecular domain due to their unsatisfactory fine-tuning performance. However, with ICMA, Mistral-7B could easily demonstrate its superior in-context learning and reasoning capabilities and perform its advantage of parameters. As a result, ICMA(Mistral-7B) achieves the best performance across all the models, obtaining 0.581 BLEU-4 and 0.661 METEOR scores on the ChEBI-20 dataset, which is even better than the domain-specific pre-trained Galactica-125M.
Cap2Mol. Similarly, as shown in Table 3, compared to their original foundation models with SFT, ICMA significantly boosts the molecule generation performance. Notably, ICMA (Mistral-7B) achieves state-of-the-art molecule generation performance, generating 46.0% exactly matched molecules, which nearly doubles the results of naive supervised fine-tuned Mistral-7B. Meanwhile, ICMA(Mistral-7B) also achieves superior performance on the three molecule fingerprints scores and obtains the best validity score, illustrating better molecule generation quality with ICMA.
Additionally, experiments are also conducted on a smaller dataset, PubChem324k. As shown in Table 4, ICMA could still boost the Mol2Cap performance of LLMs like Galactica-125M and Mistral-7B under the low-data scenario. However, due to the lack of similar examples, Table 5 demonstrates that the molecule generation capability can be influenced when context examples are not sufficient to support LLMs in making a correct decision.
[…]
Galactica-125M 0.333 0.265 0.465 0.322 0.417 0.406
ICMA(Galactica-125M)2,1024 0.383 0.311 0.479 0.341 0.428 0.438
Mistral-7B 0.361 0.288 0.471 0.325 0.419 0.421
ICMA(Mistral-7B)2,2048ß 0.386 0.314 0.483 0.345 0.431 0.443
Table 4: Mol2Cap results on PubChem324k dataset (Best, Second Best). Here, Galactica-125M and Mistral-7B demonstrate vanilla supervised fine-tuned results, while ICMA(Galactica-125M) and ICMA(Mistral-7B) illustrate the ICMA results.
[…]
Galactica-125M 0.485 0.031 62.08 0.681 0.510 0.403 0.835
ICMA(Galactica-125M)1,1024 0.569 0.035 52.75 0.719 0.579 0.652 0.825
Mistral-7B 0.438 0.082 74.16 0.731 0.577 0.472 0.866
ICMA(Mistral-7B)1,2048 0.450 0.065 77.01 0.764 0.624 0.504 0.891
Table 5: Cap2Mol results on PubChem324k dataset (Best, Second Best). Here, Galactica-125M and Mistral-7B demonstrate vanilla supervised fine-tuned results, while ICMA(Galactica-125M) and ICMA(Mistral-7B) illustrate the ICMA results.
4.3 Study of Retrieval Algorithms
[…]
Random 0.573 0.496 0.634 0.484 0.572 0.596
Morgan FTS 0.599 0.526 0.650 0.508 0.590 0.626
Mole-BERT 0.618 0.545 0.662 0.521 0.602 0.637
Table 6: Performance comparison of different retrieval algorithms for Mol2Cap task (Best, Second Best). The backbone is ICMA(Galactica-125M)2,1024.
[…]
Random 0.802 0.286 25.29 0.867 0.763 0.699 0.923
SBERT 0.795 0.245 26.05 0.868 0.765 0.691 0.926
BM25 0.818 0.293 24.20 0.881 0.784 0.717 0.927
Table 7: Performance comparison of different retrieval algorithms for Cap2Mol task (Best, Second Best). The backbone is ICMA(Galactica-125M)2,1024.
We study the influence of retrieval algorithms to illustrate the importance of retrieval qualities. For molecule retrieval, we compare the new proposed Molecule Graph Retrieval using Mole-BERT with random retrieval and the Morgan FTS retrieval. As shown in Table 6, Molecule Graph Retrieval using Mole-BERT achieves the best results on all of the metrics, proving its superiority in molecule retrieval. For the caption retrieval, we compare BM25 Caption Retrieval with random retrieval and SBERT retrieval under the framework of ICMA. As depicted in Table 7, BM25 Caption Retrieval illustrates its excellent performance among the three caption retrieval methods. Notably, SBERT retrieval achieves a worse score than random retrieval because it loses detailed information during the compression, further proving the importance of details in caption retrieval.
4.4 Study of Context Settings
Figure 3: The model performance with the change of context settings, including the number of examples (i.e., n_shot) and cutoff length. Mol2Cap Results (Left) and Cap2Mol Results (Right).
In ICMA, the context settings, including the example number and cutoff length, are also important to its performance. The increase of context example number will also drastically add to the context length. However, if the context length is larger than the cutoff length, then the training will be insufficient because most of the inputs are cropped, and the information is lost. In this case, we want to make sure that most of the context examples fit in the cutoff length. Considering that the input length limitation of Galactica-125M is 2048, we test the cutoff length within the range of {512, 1024, 1536, 2048} and the example number from 1 to 4 for analysis. As illustrated in Figure 3, when the example number increases, the performance becomes worse if the cutoff length is too short. For example, when the cutoff length is set to 512, the context is not complete for LLMs to consistently learn valuable knowledge from it, leading to worse results. However, when the cutoff length is long enough, the performance mainly becomes better as the increased example number provides more information to help LLMs make accurate predictions.
4.5 Study of Scaling Law
Figure 4: The scaling law of ICMA. Three models with different levels of parameters are selected, including Galactica-125M, Galactica-1.3B, and Mistral-7B. Mol2Cap Results (Left) and Cap2Mol Results (Right).
As ICMA creates a new paradigm for adapting powerful LLMs with billion parameters to the molecule-caption translation task, it is interesting to study the relationship between the performance and model scales. In this case, we select LLMs with different scales of parameters, ranging from 125 million to 7 billion, to study the scaling law of ICMA. Specifically, we select Galactica-125M, Galactica-1.3B, and Mistral-7B for demonstration.
As illustrated in Figure 4, with the increase of model scale, the performance of ICMA also improves. Especially in Mol2Cap task, the METEOR score has been drastically improved when the model parameters scale from 1.3 billion to 7 billion. The improvement of molecule understanding significantly benefits from the language generation capability as the model scale reaches a certain level. Meanwhile, for Cap2Mol task, the accuracy for predicting exactly matched molecules demonstrates a slightly different pattern from the Mol2Cap task. The model’s performance in generating exact-matched molecules seems to have a log-linear relationship with the model scales, showing that larger language models could explore the molecular space better owing to their more powerful capabilities. Overall, ICMA allows general LLMs with billion parameters to achieve similar or better performance to these domain-specific pre-trained LLMs in the molecule-caption translation task, showing that LLMs are inherently in-context molecule learners.
4.6 Ablation Study
[…]
ICMA 0.618 0.545 0.662(4) 0.521(1) 0.602(3) 0.637(3)
w/o Random Walk 0.603 0.534 0.661(9) 0.523(3) 0.602(1) 0.638(1)
w/o Sequence Reverse 0.597 0.528 0.661(8) 0.522(5) 0.601(8) 0.638(0)
Table 8: Ablating components of Post-retrieval Re-ranking for Mol2Cap task (Best, Second Best). The backbone is ICMA(Galactica-125M)2,1024.
[…]
ICMA 0.818 0.293 24.20 0.881 0.784 0.717 0.927
w/o Random Walk 0.771 0.103 28.99 0.838 0.749 0.654 0.925
w/o Sequence Reverse 0.807 0.279 24.65 0.876 0.782 0.713 0.929
Table 9: Ablating components of Post-retrieval Re-ranking for Cap2Mol task (Best, Second Best). The backbone is ICMA(Galactica-125M)2,1024.
Lastly, the ablation study is also conducted to illustrate how the Post-retrieval Re-ranking components affect the predictions. We conduct ablation experiments for the Mol2Cap and Cap2Mol sub-tasks by deactivating the Random Walk and Sequence Reversal. As shown in Table 8 and 9, without Random Walk or Sequence Reversal, the performance of ICMA drops in both tasks, proving the effectiveness of Post-retrieval Re-ranking.
5 Conclusion
In this work, we propose In-Context Molecule Adaptation (ICMA), as a new paradigm for adapting LLMs to the molecule-caption translation task. Instead of domain-specific pre-training and fine-tuning, ICMA enables LLMs to utilize their in-context learning capability to learn molecule-text alignment via In-Context Molecule Tuning, which significantly improves the performance of LLMs in the molecule-caption translation task, demonstrating that LLMs are inherently in-context molecule learners. More importantly, our study provides a viable framework for deploying advanced LLMs with billion-level parameters in the scientific field.
6 Limitations
The limitation of this work primarily lies in the following aspects:
  • Firstly, we limited the retrieval range to the training split of the dataset, which might be insufficient, given the vastness of the molecular space.
  • Secondly, this work is limited to the molecule-caption translation task. However, it is worth noting that there are various molecule-related downstream tasks, such as molecule property prediction, which is also important to verify the generalization capability of ICMA.
  • Lastly, due to the limitation of hardware, we were unable to investigate the potential of language models with more than 7 billion parameters with ICMA. It would be interesting to explore the capabilities of larger, more sophisticated language models under the framework of ICMA.
7 Acknowledgement
We thank any anonymous reviewers for their valuable suggestions. We also thank ChatGLM-4111https://chatglm.cn and Github Copilot [^2] https://github.com/features/copilot for assistance in the paper revision and code writing.
Appendix A Appendix
A.1 Hyper Parameters
Table 10 illustrates all the hyperparameters and their corresponding value ranges. The first part is about the general training settings and ICMA settings. Notably, if not specifically stated, the default cutoff length is set to 1024, and the n_shot (i.e., number of context examples) is set to 2. Next, Lora and acceleration settings are listed. Lastly, we also provide the hyperparameters for controlling the generation.
@@ A.1 Hyper Parameters
n_shot 1,2,3,4
@@
Table 10: Hyper-Parameters
A.2 Context Templates
Next, we introduce the context templates adopted to form the input of LLMs. Generally, the context templates are much simpler than those in MolReGPT because LLMs are fine-tuned in ICMA, which makes it easier for output-format control. In this case, the format control prompts could be omitted. At the same time, we still add proper prompts before generation to motivate LLMs to learn from the context examples, especially identifying the similarities and differences within these context examples.
Cap2Mol ICMA Prompt  Generate a caption for the molecule: {PlaceHolder}
Caption: {PlaceHolder}
Based on the above examples, analyse the similarities and differences between the examples and finally generate a caption for the molecule: {PlaceHolder}.
Caption: {For Generation}
Mol2Cap ICMA Prompt  Generate a molecule for the caption: {PlaceHolder}
Molecule:{PlaceHolder}
Based on the above examples, analyse the similarities and differences between the examples and finally generate a molecule for the caption: {PlaceHolder}.
Molecule: {For Generation}
A.3 Further Discussion
In this section, we conduct some further experiments to provide extra insights and support our method. First, we test and compare the performances of different backbone LLMs with and without ICMA. Then, we compare ICMA with some other baseline models.
A.3.1 Results with More Backbone LLMs

### v3

@@ III-B2 Sequence Reversal
III-C In-Context Molecule Tuning
As shown in Figure 1, similar molecules typically share similar structures and chemical properties. Building on this principle, MolReGPT [10] has demonstrated the effectiveness of in-context learning, which aims to prompt Large Language Models (LLMs) to learn from similar examples without the need for domain-specific pre-training and fine-tuning. However, MolReGPT heavily relies on the reasoning and in-context learning capabilities of LLMs, resulting in poor performance with relatively small language models. To address the deficits , we propose In-Context Molecule Tuning to fine-tune the parameters of LLMs, enabling them to learn from context examples and reason on the current input. Notably, In-Context Molecule Tuning can be easily adapted to any LLM, allowing even smaller language models to unlock their in-context molecule learning capability by learning the differences and similarities between molecule-caption pairs in the context.
@@ IV Experiments
In this section, we aim to evaluate the effectiveness of ICMA. Firstly, we introduce the experimental settings. Then, we compare ICMA with the selected baselines on the ChEBI-20 dataset and further test ICMA on a smaller dataset, PubChem324k. Meanwhile, we also comprehensively study the factors that will affect the performance of ICMA, including retrieval algorithms, context settings, model scales, backbone models, complexity of data samples, and the Random Walk sampling strategy. After that, an ablation study is conducted to justify the design of Post-retrieval Re-ranking components. What’s more, we compare ICMA with models that require extra-domain alignment stages to demonstrate the efficiency and effectiveness of ICMA. Furthermore, we verify ICMA on molecule property prediction tasks, illustrating the generalization capability of ICMA to other molecule-related tasks. Finally, we conduct a case study to provide more details about our method.
IV-A Experimental Settings
@@
We will first detail our experiment settings. All the hyper-parameters are illustrated in Table I. If not specifically stated, the cutoff length that truncates the inputs for LLMs to process is set to 1024, and n_shot is set to 2 to control the variables. Notably, for LLMs with over 1 billion parameters, we apply LoRA [28] to save the GPU memory and accelerate computation. Otherwise, we fine-tune the full model of LLMs. For the dataset, we apply two different molecule-caption translation datasets, ChEBI-20 [7] and PubChem324k [9]. The details of the datasets are shown in Table II.
TABLE I: Hyper-Parameters
@@
refined examples ($n$) 1,2,3,4
rough examples ($N$) 10
maximum skip probability ($p_{max}$) 9
@@
TABLE II: Details of the datasets, ChEBI-20 and PubChem324k. For PubChem324k, we follow the split in MolCA [9], while ignoring the Pretrain fold.
Dataset Train Validation Test
ChEBI-20 26,407 3,001 3,000
PubChem324k 12,000 1,000 2,000
For comparison, we first select two different foundation LLMs as the backbones of ICMA, namely Galactica-125M [29] and Mistral-7B-instruct-v0.2 [30]. The former is a representative and smaller LLM that has been pre-trained on unstructured scientific corpora, which have been aware of the molecular knowledge from the pre-training stage, while the latter is a general LLM with 7 billion parameters, whose capabilities are comparable to GPT-3.5-turbo. To study the model agnosticism, we include two more backbone LLMs (Galactica-1.3B and Llama-2-7b-chat-hf) in Section IV-F. Notably, in the factor analysis and ablation study parts, we mainly select Galactica-125M for experiments to reduce the computational costs. To better present our results, considering baseline models, we first select MolT5-base, MolT5-large [5] and MolReGPT (GPT-3.5-turbo and GPT-4-0314) [10] for comparison on the ChEBI-20 dataset and the case study. We also include the comparison and discussion with previous additional SOTA methods, namely BioT5 [12], MolXPT [11], and MolCA [9], in Section IV-J.
IV-B Performance Comparison
TABLE III: Mol2Cap results on ChEBI-20 dataset (Best, Second Best). Here, the results of MolT5-base and MolT5-large are domain-specific pre-training & fine-tuning results [5], while MolReGPT(GPT-3.5-turbo) and MolReGPT(GPT-4-0314) are prompting & in-context learning results [10]. More importantly, Galactica-125M and Mistral-7B demonstrate naive supervised fine-tuned results, while ICMA(Galactica-125M) and ICMA(Mistral-7B) illustrate the ICMA results.
[…]
MolT5-base [5] 0.540 0.457 0.634 0.485 0.578 0.569
MolReGPT (GPT-3.5-turbo) 0.565 0.482 0.623 0.450 0.543 0.585
MolT5-large [5] 0.594 0.508 0.654 0.510 0.594 0.614
MolReGPT (GPT-4-0314) 0.607 0.525 0.634 0.476 0.562 0.610
Galactica-125M 0.585 0.501 0.630 0.474 0.568 0.591
ICMA(Galactica-125M)2,2048 0.636 0.565 0.674 0.536 0.615 0.648
Mistral-7B 0.566 0.478 0.614 0.449 0.547 0.572
ICMA(Mistral-7B)2,2048 0.651 0.581 0.686 0.550 0.625 0.661
TABLE IV: Cap2Mol results on ChEBI-20 dataset (Best, Second Best). Here, the results of MolT5-base and MolT5-large are domain-specific pre-training & fine-tuning results [5], while MolReGPT(GPT-3.5-turbo) and MolReGPT(GPT-4-0314) are prompting & in-context learning results [10]. More importantly, Galactica-125M and Mistral-7B demonstrate naive supervised fine-tuned results, while ICMA(Galactica-125M) and ICMA(Mistral-7B) illustrate the ICMA results.
[…]
MolT5-base [5] 0.769 0.081 24.458 0.721 0.588 0.529 0.772
MolReGPT(GPT-3.5-turbo) 0.790 0.139 24.91 0.847 0.708 0.624 0.887
MolT5-large [5] 0.854 0.311 16.071 0.834 0.746 0.684 0.905
MolReGPT(GPT-4-0314) 0.857 0.280 17.14 0.903 0.805 0.739 0.899
Galactica-125M 0.781 0.173 26.34 0.836 0.708 0.631 0.916
ICMA(Galactica-125M)4,2048 0.843 0.391 17.71 0.897 0.812 0.753 0.941
Mistral-7B 0.767 0.234 27.39 0.852 0.718 0.649 0.918
ICMA(Mistral-7B)4,2048 0.855 0.460 18.73 0.916 0.837 0.789 0.958
Mol2Cap. As illustrated in Table III, Galactica-125M with SFT has already shown competitive performance to previous baselines due to its pre-training on scientific corpora. However, ICMA could still improve the performance of Galactica-125M by 12.8% and 8.3% considering the BLEU-4 and ROUGE-L scores on the ChEBI-20 dataset. With only 125 million parameters, ICMA(Galactica-125M) can beat MolT5-large, which owns more than 780 million parameters. Meanwhile, the general LLM, Mistral-7B with naive SFT, only achieves a performance that is slightly better than MolT5-base, despite the fact that Mistral-7B is 70 times larger than MolT5-base. This outcome is not surprising because Mistral-7B is not specifically designed or pre-trained for biomolecular purposes. It also reveals that although general LLMs have illustrated powerful capabilities with billions of parameters, few works have adapted them to the biomolecular domain due to their unsatisfactory fine-tuning performance. However, with ICMA, Mistral-7B could easily demonstrate its superior in-context learning and reasoning capabilities and perform its advantage of parameters. As a result, ICMA(Mistral-7B) achieves the best performance across all the models, obtaining 0.581 BLEU-4 and 0.661 METEOR scores on the ChEBI-20 dataset, which is even better than the domain-specific pre-trained Galactica-125M.
Cap2Mol. Similarly, as shown in Table IV, compared to their original foundation models with SFT, ICMA significantly boosts the molecule generation performance. Notably, ICMA(Mistral-7B) achieves state-of-the-art molecule generation performance, generating 46.0% exactly matched molecules, which nearly doubles the results of naive supervised fine-tuned Mistral-7B. Although both ICMA(Galactica-125M) and ICMA(Mistral-7B) achieve higher Levenshtein scores, due to the characteristic of the metric, a lower Levenshtein score does not mean better generation quality. For example, the Levenshtein score of $CO\rightarrow CCOC$ is 2, and the Levenshtein score of $CCC\rightarrow CCOC$ is just 1. However, the former molecule is considered more similar to the target molecule as they share the same functional group. Instead, if we look at the molecule fingerprints similarity scores, ICMA(Mistral-7B) could achieve superior results and obtain the best validity score, showing better molecule generation quality.
Additionally, experiments are also conducted on a smaller dataset, PubChem324k. As shown in Table V and VI, ICMA could still boost the Mol2Cap performance of LLMs like Galactica-125M and Mistral-7B, which further proves the generalization of ICMA in the molecule-caption translation task.
TABLE V: Mol2Cap results on PubChem324k dataset (Best, Second Best). Here, Galactica-125M and Mistral-7B demonstrate naive supervised fine-tuned results, while ICMA(Galactica-125M) and ICMA(Mistral-7B) illustrate the ICMA results.
[…]
Galactica-125M 0.333 0.265 0.465 0.322 0.417 0.406
ICMA(Galactica-125M)4,2048 0.411 0.338 0.497 0.359 0.446 0.457
Mistral-7B 0.361 0.288 0.471 0.325 0.419 0.421
ICMA(Mistral-7B)4,2048 0.416 0.345 0.505 0.367 0.453 0.464
TABLE VI: Cap2Mol results on PubChem324k dataset (Best, Second Best). Here, Galactica-125M and Mistral-7B demonstrate naive supervised fine-tuned results, while ICMA(Galactica-125M) and ICMA(Mistral-7B) illustrate the ICMA results.
[…]
Galactica-125M 0.485 0.031 62.08 0.681 0.510 0.403 0.835
ICMA(Galactica-125M)4,2048 0.600 0.098 49.00 0.764 0.632 0.523 0.909
Mistral-7B 0.438 0.082 74.16 0.731 0.577 0.472 0.866
ICMA(Mistral-7B)4,2048 0.526 0.163 62.25 0.799 0.678 0.573 0.935
IV-C Study of Retrieval Algorithms
TABLE VII: Performance comparison of different retrieval algorithms for Mol2Cap task (Best, Second Best). The backbone is ICMA(Galactica-125M)2,1024.
[…]
Random 0.611 0.533 0.648 0.500 0.587 0.613
Morgan FTS 0.627 0.552 0.661 0.517 0.600 0.633
Mole-BERT 0.641 0.568 0.671 0.531 0.611 0.645
TABLE VIII: Performance comparison of different retrieval algorithms for Cap2Mol task (Best, Second Best). The backbone is ICMA(Galactica-125M)2,1024.
[…]
Random 0.823 0.299 20.24 0.873 0.772 0.709 0.928
SBERT 0.828 0.322 19.91 0.880 0.786 0.720 0.929
BM25 0.842 0.355 17.99 0.890 0.805 0.741 0.935
We also study the influence of retrieval algorithms to illustrate the importance of retrieval qualities. For molecule retrieval, we compare the new proposed Molecule Graph Retrieval using Mole-BERT with random retrieval and the Morgan FTS retrieval. As shown in Table VII, Mole-BERT achieves the best results on all of the metrics, proving its superiority in molecule retrieval. For the caption retrieval, we compare BM25 Caption Retrieval with random retrieval and SBERT retrieval under the framework of ICMA. As depicted in Table VIII, BM25 Caption Retrieval illustrates its excellent performance among the three caption retrieval methods.
IV-D Study of Context Settings
Figure 3: The model performance with the change of context settings, including the number of refined examples (i.e., $n$) and cutoff length. Mol2Cap Results (Left) and Cap2Mol Results (Right).
In ICMA, the context settings, including the example number and cutoff length, are also important to its performance. The increase in example number will also drastically require longer context length. If the context length is longer than the cutoff length, then the training will be insufficient because most of the inputs are cropped, and the information is lost. Meanwhile, during the inference stage, the context length also influences the information that LLMs can take. In this case, we want to make sure that most of the context examples fit in the cutoff length. Considering that the input length limitation of Galactica-125M is 2048, and the model series has no length extrapolating capability, we test the cutoff length within the range of {512, 1024, 1536, 2048} and the example number from 1 to 4 for analysis. As illustrated in Figure 3, when the example number increases, the performance becomes worse if the cutoff length is too short. Notably, when the cutoff length is set to 512, the context is not complete for LLMs to consistently learn valuable knowledge from it, leading to worse results. However, when the cutoff length is long enough, the performance mainly becomes better as the increased example number provides more information to help LLMs make accurate predictions. This phenomenon is more obvious in the Cap2Mol task, while in the Cap2Mol task, the impact of context settings is not significant, as there is a balance between the extra information and noises with the increase of cutoff length and context example numbers.
IV-E Study of Scaling Law
Figure 4: The scaling law of ICMA. Three models with different levels of parameters are selected, including the Galactica series (Blue) and the Llama-3 series (Green). Mol2Cap Results (Left) and Cap2Mol Results (Right).
As ICMA creates a new paradigm for adapting powerful LLMs with billion parameters to the molecule-caption translation task, it is interesting to study the relationship between the performance and model scales. In this case, we select LLMs with different scales of parameters, ranging from 125 million to 8 billion, to study the scaling law of ICMA. Specifically, we select the Galactica series and Llama-3 series for demonstration.
As illustrated in Figure 4, with the increase of model scale, the performance of ICMA for the Llama-3 series improves in both the Mol2Cap and Cap2Mol tasks. The improvements in molecule understanding and generation capabilities significantly benefit from the increased model scales. In the case of the Galactica series, when the model size is increased from 1.3 billion to 6.7 billion parameters, a consistent pattern is observed in the Cap2Mol task. However, the performance on the Mol2Cap task exhibits a notably different trend, with minimal improvement observed. This discrepancy could indicate underlying deficiencies in the model’s natural language understanding and generation capabilities, which caps Galactica‘s performance in the Mol2Cap task.
IV-F Study of Model Agnosticism